Diffusion Morning Report 08/30/05

Ted Berg tberg at snf.stanford.edu
Tue Aug 30 07:03:17 PDT 2005


Good Morning All,
    Looking good this morning. all of the LP tubes are back up with the 
exception of TEOS which needs a test. Tylan 2 had a temp problem-
repaired TC connection.  New furnace install is moving along.  Epi looks 
like it has a small leak somewhere-looking. Everything else looks
OK for now.

  	Date: 	8/30/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	Repaired TC connection 	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Down 	Needs test
	4.67/30
Tylan 9-Poly 	Up 	  	20.99/30
Tylan 10-Nitride 	Up 	  	11.85/30
Tylan 11-BPSG-LTO 	Up 	  	8.45/30
Tylan 12-FGA 	Up 	  	 
  	  	  	 
  	  	  	 
  	  	  	 
  	  	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	Up 	  	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	.
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	B on line 700 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050830/fe303e6c/attachment.html>


More information about the diffusion mailing list